A photocopier increases the sides of a square in the ratio 2:3. What percentage increase is this?
Answer by Theo(13342)

let a = the length of the side of the original square.
let b = the length of the side of the increased square.

you get the ratio a/b = 2/3.

solve for b to get b = 3/2 * a.

since 3/2 = 1.5, then you get b = 1.5 * a

b is 1.5 * the size of a.

the increase is calculated by subtracting a from b.

b - a is equivalent to 1.5 * a - a because b = 1.5 * a.

1.5 * a - a becomes equal to .5 * a.

the increase is .5 * a which is equivalent to 50% * a.

a is increased by 50% of a to get b which is equal to 150% of a.

for example:

assuming a is equal to 10.

it is increased by 50% of 10 which means it is increased by 5.

the new side length is 10 + 5 = 15.

the new side length is equal to 150% of 10 which is equal to 1.5 * 10 which is equal to 15.